OpenStudy (anonymous):

Express the radical in simplest form: /121a^3b^2 the "/" = sqrt i dunno how to type it :/

OpenStudy (anonymous):

help please :(

OpenStudy (dumbcow):

\[\sqrt{121a^{3}b^{2}}\] right?

OpenStudy (anonymous):

yes:)

OpenStudy (dumbcow):

ok use this important property \[\sqrt{x^{2}} = x\] so try to split everything into groups of 2 also 121 = 11^2 --> 11*11 = 121 \[\rightarrow \sqrt{11^{2}a^{2}ab^{2}}\] i turned a^3 into a^2*a now everything that is squared can be taken out of radical \[\rightarrow 11ab \sqrt{a}\]

OpenStudy (anonymous):

Would it be 11ab/a

OpenStudy (dumbcow):

nope just as i wrote it \[11ab \sqrt{a}\]

OpenStudy (anonymous):

lol i dunno how to do the sqrt symbol :P But is that the right answer?

OpenStudy (dumbcow):

yes it is haha go to equation button and type sqrt{ x }

OpenStudy (anonymous):

thank you! :)